If the product of two numbers is 56 and the sum is 15,what are the two numbers?
this is a fifth grade question though.

Answer by stanbon(75887)

If the product of two numbers is 56 and the sum is 15,what are the two numbers?
------
Equations:
xy = 56
x + y = 15
------
I would guess that a 5th grader would ask him or herself
what numbers give 56 when they are multiplied.
That would be 1/56 ; 2/28 ; 4/14 ; 7/8
----
Only 7/8 add up to 15.
So x = 7 and y = 8 or x = 8 and y = 7
=========================================
If the problem is done with algebra it looks like the following:
x = 15-x
Substitute for "x" and solve for "y":
(15-x)x = 56
15x - x^2 = 56
x^2 - 15x + 56 = 0
Factor:
(x-7)(x-8) = 0
x = 7 or x = 8
---
If x = 7, y = 8
If x = 8, y = 7
========================
Maybe your 5th grader knows that much algebra.
=========================
Cheers,
Stan H.
========================